OpenStudy (anonymous):

without applying l'hospital rule what is the answer of lim x->inf e^x / x^2, how come is it inf as the answer of lim x->inf e^x is inf & the answer of lim x->inf 1/x^2 is 0. and the answer of lim x->inf e^x /x^2 will be inf/0, all I want to ask is that can we divide inf by zero ?

OpenStudy (phi):

as x --> infinity, you get e^infty/infty^2 but e^x grows faster than x^2, so the expression goes to infinity.

OpenStudy (anonymous):

that's some scary dp bro..

OpenStudy (anonymous):

phi; how come lim x->inf 1/x^2 is inf? as http://tutorial.math.lamar.edu/Classes/CalcI/LimitsAtInfinityI.aspx it says that it is zero

OpenStudy (phi):

you were asked about \[ \frac{e^x}{x^2} \]

OpenStudy (phi):

which is a ratio of two values that approach infinity

OpenStudy (phi):

or you can think of it as \[ e^x \cdot \frac{1}{x^2} \] which you might view as infinity times zero in the limit.

OpenStudy (anonymous):

yeah so if we do that like lim x->inf e^x /x^2 = limx-.inf e^x * limx->inf 1/x^2 = inf *0 ohhhh my misstake my question is now that if what is the answer of inf * 0

OpenStudy (anonymous):

opppsss sorry i was dividing inf by zero , is inf * 0 = inf ?

OpenStudy (phi):

no inf*0 is undefined.

OpenStudy (phi):

to answer the question, treat it as a ratio of two quantities that both go to infinity.

OpenStudy (phi):

Then notice that an exponential "grows faster" than x^2, so in the limit, it dominates.

OpenStudy (anonymous):

OKKKKKKKKKK thanks

OpenStudy (phi):

we could use the series definition of e^x to show this.

OpenStudy (phi):

e^x = 1+x + x^2/2! + x^3/3! + divide each term by x^2 and take the limit.

OpenStudy (amistre64):

just for the practice :) and prolly wrong to begin with \[\ lim\frac{e^x}{x^2}\frac{(x-1)^2}{e^{x-1}}\] \[\ lim\frac{e}{x^2}\frac{(x-1)^2}{1}\] \[\ lim\frac{e}{x^2}\frac{x^2-2x+1}{1}\] \[\ lim\ e\left(\frac{x^2}{x^2}-\frac{2x}{x^2}+\frac{1}{x^2}\right)\] \[\ lim\ e\] hmm

OpenStudy (anonymous):

thanks amistre

OpenStudy (amistre64):

yep, i just cant recall what the result is spose to tell us :)